A spring of natura length 1.5m is extended 0.0050 by a force of 0.8n . What will its length be when the applied force is 3.2n
1
Expert's answer
2019-05-09T11:35:40-0400

Hooke's law states

"F=kx"

So

"\frac{F_1}{F_2}=\frac{x_1}{x_2}"

"x_2=\frac{F_2}{F_1}x_1"

"=\frac{3.2\:\rm{N}}{0.8\:\rm{N}}\times 0.005\:\rm{m}=0.02\:\rm{m}"

The spring length

"l_2=l_0+x_2"

"=1.5\:\rm{m}+0.02\:\rm{m}=1.52\:\rm{m}"